A. Find the mean (𝝁𝑿), variance (𝝈 𝑿

𝟐) and standard deviation (πˆπ‘Ώ) of the sampling distribution of the means given the sample size, and the means and variances of the population. Sample size n were randomly selected from the population.


1. 𝑛 = 3

πœ‡ = 6

𝜎² = 2.4

Mean:


Variance:


Standard Deviation:


2. 𝑛 = 25

πœ‡ = 20

𝜎² = 5.5

Mean:


Variance:


Standard Deviation:


1
Expert's answer
2022-04-05T07:09:35-0400

We'll use the properties of sampling distributions of sample means.


1. Mean:

"\mu_{\bar x} =\mu=6."

Variance:

"\sigma^2_{\bar x}=\cfrac{\sigma^2}{n}=\cfrac{2.4}{3}=0.8."

Standard deviation:

"\sigma_{\bar x}=\cfrac{\sigma}{\sqrt n}=\sqrt{\sigma^2_{\bar x}}=\sqrt{0.8}=0.89."


2. Mean:

"\mu_{\bar x} =\mu=20."

Variance:

"\sigma^2_{\bar x}=\cfrac{\sigma^2}{n}=\cfrac{5.5}{25}=0.22."

Standard deviation:

"\sigma_{\bar x}=\cfrac{\sigma}{\sqrt n}=\sqrt{\sigma^2_{\bar x}}=\sqrt{0.22}=0.47."